# From Concept to Measurement: A Survey of How the Blockchain Trilemma Can Be Analyzed

### v1

@@ I Introduction
Selecting suitable constructs and metrics for analyzing the blockchain trilemma is challenging due to the insufficient development of its theoretical foundations. Without such theoretical foundations, the suitability of constructs and their operationalization through metrics for investigations of the blockchain trilemma can hardly be justified. To tackle the lack of such a theoretical foundation, extant research presents mappings of the blockchain trilemma to established theories in distributed systems, such as the CAP theorem [3, 15, 16]. However, such mappings often do not cover all subconcepts or intermingle constructs of different subconcepts. Although such adjacent theoretical foundations are valuable, they only reflect parts of the blockchain trilemma, which hinders thorough suitability assessments of constructs for analyzing the blockchain trilemma. Due to inconsistencies in construct usage and a lack of a well-defined link between constructs and the blockchain trilemma subconcepts, practitioners must rely on individual discretion to identify constructs suitable for Pareto-optimal configurations that meet non-functional requirements. To assist practitioners in selecting appropriate constructs and associated metrics to analyze the blockchain trilemma subconcepts, we pose the following research question: What constructs and associated metrics are suitable to quantify the blockchain trilemma subconcepts?
We conducted a systematic literature search [17] to identify relevant publications discussing constructs and their associated metrics used to analyze subconcepts of the blockchain trilemma. Next, we used open, axial, and selective coding [18, 19] to analyze the selected publications. In open coding, we extracted constructs and their associated metrics, which serve to operationalize the subconcepts of the blockchain trilemma across different blockchain systems. In axial coding, we related the constructs to specific blockchain trilemma subconcepts (i.e., DoD, scalability, and security). Finally, through selective coding, we refined the assignment of constructs to blockchain trilemma subconcepts to ensure theoretical consistency and coherence. Based on the literature review, we developed an overview of analysis approaches that use the identified constructs and metrics to investigate the blockchain trilemma.
@@
The primary purpose of this work is to enhance understanding of the constructs and associated metrics used to analyze blockchain trilemma and its subconcepts. In particular, this work has three main contributions. First, by explaining common constructs and their operationalization through metrics used to quantify DoD, scalability, and security, their applicability, interpretability, and limitations, we offer a theoretical foundation for more targeted analyses of blockchain trilemma subconcepts. Second, by explaining those metrics, including their input variables, we support data collection in benchmarks. For example, we clarify blockchain system characteristics that need to be monitored to feed input variables of relevant metrics used in benchmarks. Third, by comparing analysis approaches based on the constructs and metrics used, we guide the selection of suitable approaches for investigating the blockchain trilemma.
@@ II-B Blockchain Trilemma Subconcepts and Their Interrelationships
Figure 2: Interrelationships between the blockchain trilemma, subconcepts, constructs, metrics.
@@ II-C Related Research on Measuring the Blockchain Trilemma Subconcepts
In addition to conceptual works, empirical studies propose multiple constructs and metrics to quantify the blockchain trilemma subconcepts. Focusing on DoD, scalability or security, in separation–not the entire blockchain trilemma–prior research [8, 7, 9, 55, 26] conducted experiments to better understand influences of blockchain system configurations on these subconcepts. Those works present multiple constructs for the same blockchain trilemma subconcepts. For example, changes in throughput and changes in confirmation latency due to faulty nodes were proposed to calculate fault tolerance, [8, 7] and stale block rate [56, 10] is used to operationalize the security subconcept. Similarly, throughput (i.e., transactions per second; e.g., [8, 7, 57]) and confirmation latency (e.g., [9, 58, 59, 56]) are used to estimate scalability subconcept. The same constructs are used to investigate different blockchain trilemma subconcepts. For example, availability was proposed to investigate scalability [9] and security [10]. Selecting suitable constructs remains difficult because there is insufficient justification for the validity of the proposed constructs and their operationalizations for measuring the blockchain trilemma subconcepts and how to interpret the constructs.
@@
In summary, prior research proposes useful constructs and corresponding operationalization through metrics to quantify the blockchain trilemma subconcepts in blockchain systems with different designs. However, justification for the suitability of proposed constructs is often lacking, and literature on operationalized constructs to measure the blockchain trilemma subconcepts is scattered across various sources. This leads to difficulty in selecting suitable constructs and metrics. To tackle this issue, the suitability of constructs and associated metrics for measuring the blockchain trilemma subconcepts needs to be better understood.
@@ III Methods
We developed a set of constructs and associated metrics to operationalize the subordinate concepts of the blockchain trilemma in two steps. First, we performed a systematic literature search [17] to compile an extensive set of relevant publications on the blockchain trilemma. Second, we analyzed the collected literature using open coding, axial coding, and selective coding [18, 19] to extract the constructs and associated metrics used to measure the blockchain trilemma subconcepts. The following details these two steps.
@@ III-A Literature Search
We conducted a systematic literature search [17] to identify relevant publications presenting constructs and associated metrics to analyze the blockchain trilemma subconcepts. First, we defined the inclusion criteria English language, level of detail, peer-reviewed, topic fit, and uniqueness to systematically assess the relevance of publications (see Table I).
@@
We used the search string: (“benchmarking” AND “blockchain trilemma”) to compile a set of publications on the blockchain trilemma via ACM Digital Library, IEEEXplore, ScienceDirect, and Scopus on March 26, 2024. We selected this search string based on preliminary reviews of existing literature and domain-specific terminology. The literature search yielded 1,814 potentially relevant publications: 1,258 publications from ACM Digital Library, 546 publications from IEEEXplore, 7 publications from ScienceDirect database, and 3 publications from Scopus database.
Next, we assessed the 1,814 publications regarding their relevance to answering our research question using the inclusion criteria presented in Table I. We applied the inclusion criteria to each publication’s title, keywords, and abstract. We excluded all publications that do not meet all inclusion criteria from further consideration. We excluded 436 publications: 4 publications were not in English, 348 publications were not peer-reviewed, 77 publications had insufficient topic fit, and 7 publications were duplicates. The relevance assessment based on title, keyword, and abstract led us to a set of 1,378 potentially relevant publications.
@@
The full-text relevance assessment revealed a gap in the collected publications adequately addressing DoD and security. We conducted theoretical sampling [18] via Google Scholar to supplement our dataset with additional relevant publications. To ensure consistency, we used the same search string used in the initial literature search and obtained additional publications not included in the initial search. Through theoretical sampling, we obtained 12 additional publications that meet the inclusion criteria (see Table I). This led us to a final set of 155 relevant publications.
@@ III-B Literature Analysis
We used open, axial, and selective coding [18, 19] to identify constructs and metrics for analyzing the blockchain trilemma subconcepts. In parallel, we extracted analysis approaches that use extracted constructs and metrics for analyzing the blockchain trilemma subconcepts. For each analysis approach, we noted the constructs and metrics used.
In open coding, we analyzed the set of relevant publications to identify constructs and associated metrics used to measure the blockchain trilemma subconcepts. We extracted quotes from analyzed publications and assigned a name (code) to the quote that reflects a candidate construct (e.g., fault tolerance or throughput). Moreover, we developed initial definitions of each code. For the metrics, we used the corresponding output variables as codes and the metrics as definitions. We independently analyzed the publications to reduce the risk of misinterpretation and subjective bias.
After analyzing a subset of publications, we discussed the codes and corresponding definitions to identify inconsistencies and semantic overlaps of the codes. When the discussion revealed inconsistencies and semantic overlaps, we refined the coding to achieve mutual exclusiveness of the codes while maintaining collective exhaustiveness. For example, we merged robustness with fault tolerance to a single construct associated with the security subconcept. We also merged success rate and availability constructs for investigating the scalability subconcept. For the merged output variables, we kept the constructs to report different approaches to calculate the same output variables. When we reached a unanimous agreement on the refined coding, they proceeded to analyze the next subset of publications to prepare for the next discussion round until all publications were analyzed. Overall, we developed 397 initial codes in open coding. Through iterative refinement of the coding results, the final coding included 14 different codes.
In open coding, we identified issues with the explanation and representation of one construct, including flaws in the presented construct and misinterpretation of the input variables used in the associated metrics. To resolve that issue, we asked the authors for clarification. Despite multiple inquiries, the authors, however, did not respond until the date of submission, so we excluded the construct from this work.
After open coding, we used axial coding to map the operationalized constructs to the blockchain trilemma subconcepts. The mappings were based on the definitions of the blockchain trilemma subconcepts in section II-B and explanations of the purpose of the identified constructs reported in corresponding publications. We individually prepared mappings of constructs to the blockchain trilemma subconcepts. For example, we mapped throughput and confirmation latency to scalability because (1) the analyzed publications mainly reported that the constructs associated with these output variables are used for scalability analyses of blockchain systems and (2) the scope of the constructs matches the definition of scalability in section II-B2. Analogously, we assigned fault tolerance and stale block rate to security. Next, we compared the individual mappings to identify potential improvements. We identified minor inconsistencies and resolved them in discussions when a unanimous agreement was reached.
In selective coding, we assigned the coding results to the concept of the blockchain trilemma. We checked the assignment of constructs to blockchain trilemma subconcepts to ensure theoretical consistency and coherence. We identified minor inconsistencies in the definitions of the constructs and corrected them.
@@ IV Constructs, Metrics, and Analysis Approaches
This section first presents an overview of the constructs used to operationalize the blockchain trilemma subconcepts in section IV-A. We link the constructs to the blockchain trilemma subconcepts, explain the operationalization of the constructs through merits, and point out limitations of the operationalized constructs. Moreover, we offer examples of how the constructs can be used. Subsection IV-B showcases uses of the operationalized constructs in analysis approaches.
@@ IV-A1 Degree of Decentralization
$Gini$ coefficient can be used to calculate wealth distribution [11, 34, 2, 58, 12]. $Gini$ coefficient is a measure of inequalities (e.g., in terms of token ownership of validating nodes in a blockchain system) and is formalized as follows [58, 11, 34]:
@@ IV-A3 Security
The following example illustrates the use of equation 34. Imagine two failures in one year in a blockchain system, totaling 30 minutes. The $TotoalOperatingTime$ equals 525,570 minutes for a year with 365 days (525,600 minutes). Due to the two failures, using equation 27, $MTBF$ equals 262,785 minutes. Using equation 34, the reliability of the blockchain system for a year $R{(525,600)}$ is calculated as follows:

### v2

@@ I Introduction
Selecting suitable constructs and metrics for analyzing the blockchain trilemma is challenging due to the insufficient development of its theoretical foundations. Without such theoretical foundations, the suitability of constructs and their operationalization through metrics for investigations of the blockchain trilemma can hardly be justified. To tackle the lack of such a theoretical foundation, extant research presents mappings of the blockchain trilemma to established theories in distributed systems, such as the CAP theorem [3, 15, 16]. However, such mappings often do not cover all subconcepts or intermingle constructs of different subconcepts. Although such adjacent theoretical foundations are valuable, they only reflect parts of the blockchain trilemma, which hinders thorough suitability assessments of constructs for analyzing the blockchain trilemma. Due to inconsistencies in construct usage and a lack of a well-defined link between constructs and the blockchain trilemma subconcepts, practitioners must rely on individual discretion to identify constructs suitable for Pareto-optimal configurations that meet non-functional requirements. To assist practitioners in selecting appropriate constructs and associated metrics to analyze the blockchain trilemma, we pose the following research question: What constructs and associated metrics are suitable to quantify the blockchain trilemma subconcepts?
We conducted a systematic literature search [17] to identify publications that propose constructs and metrics for analyzing the subconcepts of the blockchain trilemma. Using abductive thematic analysis [18, 19, 20], we iteratively analyzed the literature to develop and refine these constructs and their associated metrics for measuring DoD, scalability, and security. The process was guided by repeated interplay between empirical findings and theoretical framing, supplemented by targeted theoretical sampling to address identified gaps and validate emerging insights. Based on this iterative abductive analysis, we developed an overview of analysis approaches that apply these constructs and metrics to investigate the blockchain trilemma.
@@
The primary purpose of this work is to enhance understanding of the constructs and associated metrics used to analyze the blockchain trilemma and its subconcepts. In particular, this work has three main contributions. First, by explaining common constructs and their operationalization through metrics used to quantify DoD, scalability, and security, their applicability, interpretability, and limitations, we offer a theoretical foundation for more targeted analyses of blockchain trilemma subconcepts. Second, by explaining those metrics, including their input variables, we support data collection in benchmarks. For example, we clarify blockchain system characteristics that need to be monitored to feed input variables of relevant metrics used in benchmarks. Third, by comparing analysis approaches based on the constructs and metrics used, we guide the selection of suitable approaches for investigating the blockchain trilemma.
@@ II-B Blockchain Trilemma Subconcepts and Their Interrelationships
Figure 2: Interrelationships between the blockchain trilemma, subconcepts, constructs, and metrics.
@@ II-C Related Research on Measuring the Blockchain Trilemma Subconcepts
In addition to conceptual works, empirical studies propose multiple constructs and metrics to quantify the blockchain trilemma subconcepts. Focusing on DoD, scalability or security, in separation–not the entire blockchain trilemma–prior research [8, 7, 9, 56, 27] conducted experiments to better understand influences of blockchain system configurations on these subconcepts. Those works present multiple constructs for the same blockchain trilemma subconcepts. For example, changes in throughput and changes in confirmation latency due to faulty nodes were proposed to calculate fault tolerance, [8, 7] and stale block rate [57, 10] is used to operationalize the security subconcept. Similarly, throughput (i.e., transactions per second; [8, 7, 58]) and confirmation latency (e.g., [9, 59, 60, 57]) are used to estimate scalability subconcept. The same constructs are used to investigate different blockchain trilemma subconcepts. For example, availability was proposed to investigate scalability [9] and security [10]. Selecting suitable constructs remains difficult because there is insufficient justification for the validity of the proposed constructs and their operationalizations for measuring the blockchain trilemma subconcepts and how to interpret the constructs.
@@
In summary, prior research proposes useful constructs and associated metrics to quantify the blockchain trilemma subconcepts in blockchain systems with different designs. However, justification for the suitability of proposed constructs is often lacking, and literature on operationalized constructs is scattered across various sources. This leads to difficulty in selecting suitable constructs and metrics. To tackle this issue, the suitability of constructs and associated metrics for measuring the blockchain trilemma subconcepts needs to be better understood.
@@ III Methods
We developed a set of constructs and associated metrics to operationalize the subordinate concepts of the blockchain trilemma in two steps. First, we conducted a systematic literature search [17] to compile an extensive set of relevant publications on the blockchain trilemma. Second, we analyzed the collected literature using abductive thematic analysis [18] to extract the constructs and associated metrics used to measure the blockchain trilemma subconcepts. The following subsections detail these two steps.
@@ III-A Literature Search
We conducted a systematic literature search [17] to identify publications that present constructs and associated metrics for analyzing blockchain trilemma subconcepts. To evaluate the relevance of publications, we applied five inclusion criteria: English language, level of detail, peer-reviewed, topic fit, and uniqueness (see Table I).
@@
We used the search string: (“benchmarking” AND “blockchain trilemma”) to compile a set of publications on the blockchain trilemma via ACM Digital Library, IEEEXplore, ScienceDirect, and Scopus on March 26, 2024. This query was informed by a preliminary review of domain-specific terminology and indexing practices. The search returned 1,814 potentially relevant publications: 1,258 from ACM Digital Library, 546 from IEEE Xplore, 7 from ScienceDirect, and 3 from Scopus.
We screened all 1,814 publications based on title, keywords, and abstract against our inclusion criteria. This step excluded 436 publications: 4 were not in English, 348 were not peer-reviewed, 77 lacked topic fit, and 7 were duplicates, leaving 1,378 potentially relevant records.
@@
During our abductive analysis (Section III-B), we observed underrepresentation of constructs related to DoD and security. To address this and enhance theoretical sufficiency, we conducted targeted theoretical sampling [19], a purposive strategy used in abductive research to refine and deepen emerging conceptual insights. We used Google Scholar and re-applied the original search string to identify studies omitted in the initial search due to indexing limitations. This supplemental search yielded 12 additional publications that met our inclusion criteria (Table I), resulting in a final corpus of 155 publications.
@@ III-B Literature Analysis
We applied abductive thematic analysis [18, 19, 20] to identify constructs and metrics associated with the blockchain trilemma subconcepts. Abductive thematic analysis combines inductive coding and deductive theorizing in an iterative process, allowing researchers to move between data and theoretical constructs to generate conceptually rich themes. This approach enables theory development that is both grounded in the literature and informed by existing conceptual frameworks [19].
We adopted the blockchain trilemma and its subconcepts (see Section II-B) as a theoretical lens. These subconcepts are inherently broad and abstract; hence, we sought to enrich and refine them through inductive engagement with the literature. Guided by abductive reasoning, we iteratively moved between data and theory, adjusting our understanding of both as patterns emerged.
We began by inductively coding passages from the 155 publications that referenced constructs or metrics relevant to the trilemma. Initial codes (e.g., fault tolerance, throughput) were derived from the data. These were continuously refined through theoretical reflection, developing a two-way relationship between emerging empirical codes and conceptual understanding.
We defined a theme as a construct that was empirically grounded and associated with at least one metric. Themes thus captured both descriptive patterns and theoretical relevance within the blockchain trilemma. To reduce redundancy, overlapping constructs were merged. For example, robustness was subsumed under fault tolerance, and availability and success rate were grouped as indicators of scalability.
To ensure reliability and reduce subjective bias, multiple researchers independently coded subsets of the literature. We resolved discrepancies through discussion, refining our thematic structure iteratively. Through cycles of coding, comparison, and theoretical integration, we converged on a stable set of 14 final themes, distilled from 397 initial codes.
One construct was excluded due to conceptual inconsistency and lack of empirical support. Despite attempts to contact the original authors for clarification, the issue remained unresolved. In line with abductive logic, which emphasizes conceptual clarity and explanatory adequacy, we excluded this construct from our results.
We then mapped each theme to one of the three trilemma subconcepts based on both patterns in the literature and alignment with conceptual definitions (Section II-B). For instance, we assigned throughput and confirmation latency to scalability, and fault tolerance and stale block rate to security. Mapping disagreements were resolved through collaborative review and consensus.
In a final review step, we assessed the internal coherence and theoretical saturation of the thematic structure. No new themes emerged during the last coding iterations, indicating thematic saturation. This confirmed that the identified constructs were both empirically grounded and conceptually robust within the blockchain trilemma framework.
@@ IV Constructs, Metrics, and Analysis Approaches
This section first presents an overview of the constructs used to operationalize the blockchain trilemma subconcepts in section IV-A. We link the constructs to the blockchain trilemma subconcepts, explain the operationalization of the constructs through merits, and point out the limitations of the operationalized constructs. Moreover, we offer examples of how the constructs can be used. Subsection IV-B showcases uses of the operationalized constructs in analysis approaches.
@@ IV-A1 Degree of Decentralization
$Gini$ coefficient can be used to calculate wealth distribution [11, 35, 2, 59, 12, 38]. $Gini$ coefficient is a measure of inequalities (e.g., in terms of token ownership of validating nodes in a blockchain system) and is formalized as follows [59, 11, 35]:
@@ IV-A3 Security
Moreover, such performance changes reflected in fault tolerance are hardly useful in benchmarks of a single blockchain system to identify its Pareto-optimal configuration: fault tolerance is, from a theoretical perspective, set with a maximum value (e.g., one-third of validating nodes). This maximum may diverge from the actual system behavior, which surely is important to be evaluated. However, to measure trade-offs between fault tolerance and constructs of DoD and scalability, different consensus mechanisms with different theoretical fault tolerance should be used.
@@
The following example illustrates the use of equation 34. Imagine two failures in one year in a blockchain system, totaling 30 minutes. The $TotalOperatingTime$ equals 525,570 minutes for a year with 365 days (525,600 minutes). Due to the two failures, using equation 27, $MTBF$ equals 262,785 minutes. Using equation 34, the reliability of the blockchain system for a year $R{(525,600)}$ is calculated as follows: